# Sub-$n^{k}$ Deterministic algorithm for minimum $k$-way cut in simple graphs

### v1

@@ 3.1 PSP-compatibility of OPT
Let $F$ be a minimum $k$-cut and let $Q$ be the partition of $V$ into connected components of GFG-F.
Assume for contradiction that $A_{\leq j-1}\nsubseteq F$, and set $S:=A_{\leq j-1}\setminus F\neq\emptyset$. Let $\widetilde{P}$ be the partition obtained from $P_{j-1}$ by gluing together those parts that are connected by edges of $S$ (equivalently: take the connected components of the graph obtained by contracting each $C\in P_{j-1}$ to a supernode and keeping exactly the superedges induced by $S$). Then $\widetilde{P}$ is a coarsening of $P_{j-1}$ whose parts are connected in $G$ (each is a union of connected $P_{j-1}$-parts linked by actual edges). Write
@@
(3) Recovering cost by undoing $d$ internal splits. Because $j$ is the first index with $\kappa(P_{j})\geq k$, we have $\kappa(P_{j-1})<k$ and thus any $k$-cut refining $P_{j-1}$ must create at least $R$ additional components inside the $P_{j-1}$-parts. If $F$ merges $d$ of these parts (so $\widetilde{P}$ has $d$ fewer), then to still reach $\geq k$ components it must create at least $R+d$ components inside the $P_{j-1}$-parts. In particular, after adding $S$ we have at least $d$ surplus components beyond $k$, so we may merge $d$ pairs of components inside $P_{j-1}$-parts and still remain a $k$-cut.
Now fix any $C\in P_{j-1}$. By ˜2.2(ii) and Theorem˜3.1, we have $\Phi(G[C])\geq\lambda_{j}$. Since any cut is a $2$-part partition, every internal cut in $G[C]$ (including singletons) has value at least $\Phi(G[C])\geq\lambda_{j}$. Therefore, each time we merge two components inside some $G[C]$ by restoring all edges between them, we save at least $\lambda_{j}$ edges from the cut set.
Perform $d$ such internal merges starting from $F^{+}$ to obtain a new cut $\widehat{F}$ with $\kappa(G-\widehat{F})\geq k$ and
@@
since $\lambda_{j-1}<\lambda_{j}$. This contradicts the optimality of $F$. Hence $S=\emptyset$ and $A_{\leq j-1}\subseteq F$.

### v2

@@ 3.1 PSP-compatibility of OPT
Let $F$ be a minimum $k$-cut and let $Q$ be the partition of $V$ into connected components of GFG-F. Assume for contradiction that $A_{\leq j-1}\nsubseteq F$, and set $S:=A_{\leq j-1}\setminus F\neq\emptyset$. Let $\widetilde{P}$ be the partition obtained from $P_{j-1}$ by gluing together those parts that are connected by edges of $S$ (equivalently: take the connected components of the graph obtained by contracting each $C\in P_{j-1}$ to a supernode and keeping exactly the superedges induced by $S$). Then $\widetilde{P}$ is a coarsening of $P_{j-1}$ whose parts are connected in $G$ (each is a union of connected $P_{j-1}$-parts linked by actual edges). Write
@@
(3) Recovering cost by undoing $d$ internal splits. By (2), we have $\kappa(G-F^{+})\geq k+d$, so $G-F^{+}$ has $d$ surplus components beyond $k$. We will restore edges inside parts of $P_{j-1}$ to reduce the number of components by $d$ while saving at least $\lambda_{j}$ edges per merge.
Fix any $C\in P_{j-1}$. Since $C$ is not split by the ideal-load recursion before level $j$, we have $\Phi(G[C])\geq\lambda_{j}$ (by ˜2.2(ii) and Theorem˜3.1). Using $\Phi(G[C])\leq\lambda(G[C])$ from ˜2.2(i), it follows that $\lambda(G[C])\geq\lambda_{j}$, and hence every non-trivial cut in $G[C]$ has size at least $\lambda_{j}$: for all $\emptyset\subsetneq U\subsetneq C$,
$|\delta_{G[C]}(U)|\geq\lambda_{j}.$
Set $F^{(0)}:=F^{+}$. For $t=0,1,\dots,d-1$: since $A_{\leq j-1}\subseteq F^{(t)}$ (we only restore edges within a single $P_{j-1}$-part), no edge remains between distinct parts of $P_{j-1}$ in $G-F^{(t)}$. If every $C\in P_{j-1}$ induced a connected subgraph in $G[C]-F^{(t)}$, then $\kappa(G-F^{(t)})\leq\kappa(P_{j-1})<k$, contradicting $\kappa(G-F^{(t)})\geq k+(d-t)>k$. Therefore, there exists some $C\in P_{j-1}$ such that $G[C]-F^{(t)}$ has at least two components. Let $U\subsetneq C$ be the vertex set of any one component of $G[C]-F^{(t)}$ and define
$X_{t}:=\delta_{G[C]}(U).$
Because $U$ is a connected component of $G[C]-F^{(t)}$, every edge of $X_{t}$ is cut by $F^{(t)}$, hence $X_{t}\subseteq F^{(t)}$. Define
$F^{(t+1)}:=F^{(t)}\setminus X_{t}.$
Restoring $X_{t}$ merges $U$ with $C\setminus U$, so $\kappa(G-F^{(t+1)})=\kappa(G-F^{(t)})-1$, and moreover $|X_{t}|=|\delta_{G[C]}(U)|\geq\lambda_{j}$, hence
$|F^{(t+1)}|\leq|F^{(t)}|-\lambda_{j}.$
After $d$ iterations let $\widehat{F}:=F^{(d)}$. Then $\kappa(G-\widehat{F})\geq k$ and
@@
using $|S|\leq d\lambda_{j-1}$ from (1) and $\lambda_{j-1}<\lambda_{j}$. This contradicts the optimality of $F$. Hence $S=\emptyset$ and $A_{\leq j-1}\subseteq F$.